### v1

@@ 4.1 Partition of $G[S]$
      Assume $\tilde{A}\cup C$ induces a balanced clustering. We set $\tilde{C}\mathrel{\mathop{:}}=C$, $\tilde{F}=\tilde{A}\cup\tilde{C}$, and claim that we have Case (2) of the lemma. As mentioned above, the existence of the flow is guaranteed by Remark 3.10. Each node $x_{a}\in X_{\tilde{A}}=X_{F}\cap A^{\prime}$ receives at most one unit of flow, and the rest receive zero.
      The fact that $A^{\prime}$ $\Omega(\phi)$-expands with respect to $\mu_{X_{F}}$ in $G^{\prime}$ means that $G^{\prime}$ is an $\Omega(\phi)$ expander with respect to $A^{\prime}\cap X_{F}=X_{\tilde{A}}$, which by definition means that $X_{\tilde{A}}$ $\Omega(\phi)$-expands in $G^{\prime}$. As $\phi=\Theta(\frac{1}{\log n})$, this means that $X_{\tilde{A}}$ $\Omega(\frac{1}{\log n})$-expands in $G^{\prime}$. This shows we have Case (2) of the Lemma.
@@ 4.2 Finding a Bottleneck Cut
As mentioned in Section 2, we want the set $F\cup B$ of inter-cluster edges and boundary edges to expand in $G[S]$. However, the existence of bottleneck-cuts may make such a partition impossible (in the cut-sparsifier case, a bottleneck cut is a cut that certifies that the boundary edges are not expanding within the cluster). Following [racke2014computing], the solution that we use in this case, is the following: As Claim 5.2 shows, in order to certify that a tree is a cut-sparsifier, we should check that all flow problems which can be routed in the tree are respected by $G$. We therefore find a “problematic” cut $(L,R=S\setminus L)$ and add $L,R$ to the tree. This forces every flow problem which is respected by $T$ to respect the cut $(L,V\setminus L)$. Specifically, we will add a cut which “blocks” the sending of flow from $X_{F}$ to $X_{B}$. As shown in Section 5, this will be enough.
Formally, the second partition is parameterized by $0<\uptau\leq 1$. We shift our focus from the graph $G[S]^{\prime}$ to the graph $G^{\prime}[S^{\prime}]$, where $S^{\prime}\mathrel{\mathop{:}}=S\cup\{x_{e}\mid e\cap S\geq 1\}$. That is, in addition to the vertices of $G[S]^{\prime}$, we also include the nodes $x_{e}$ for every boundary edge $e\in B$ and connect $x_{e}$ to $v$ where $v\in e\cap S$.
Next, we define the weighting $\mu_{\uptau}$ on the vertices of $G^{\prime}[S^{\prime}]$: the boundary split nodes are assigned a weight of $\uptau$, and the rest of the vertices nodes have weight 1.  [^7] For non split nodes, $v\in S$, $\mu_{\uptau}(v)$ does not play a role in the following discussion. As explained above, we are looking for a cut which blocks the sending of flow from $X_{F}$ to $X_{B}$. We will therefore apply the following lemma on the sets $X_{F}$ and $X_{B}$. The set of edges $Y$ will be the cut edges that “block” the flow.
Lemma 4.6 (Improved version of [racke2014computing, Lemma 3.9]).
@@
Consider the flow network where we add to $G^{\prime}$ a source $s$ connected to $X_{A}$ (with edges $(s,x_{a})$ of capacity $\mu(x_{a})$) and a sink $t$ connected to $X_{B}$ (with edges $(x_{b},t)$ of capacity $\mu(x_{b})$). It can be seen that an exact min–cut $Y$ that separates $s$ from $t$ in this network would satisfy the requirements of Lemma 4.6 with the constant $1$ instead of $2$. However, computing a min–cut is too expensive for our purposes. Instead, we use fair–cuts [LNPSsoda23], which provide the same guarantees up to constant factors and can be computed in near-linear time.
Definition 4.8 (Fair Cut, [LNPSsoda23]).
@@
Standard approximate max-flow algorithms have the unpleasant property that the flow provided can go backwards from $T$ to $S$ across the approximate min-cut $(S,T)$. Fair-cuts give a much stronger guarantee: a Fair-cut relaxes the notion of a minimum cut by requiring that every edge that crosses the cut is roughly saturated. Thus, the flow can only cross the cut from $S$ to $T$ and the capacity of every cut edge is almost fully utilized. For all of our use cases, fair-cuts give the same guarantees as min-cut, up to a constant factor.
The following theorem, due to [li2025simple], is the state of the art result for the computation of fair-cuts.
Theorem 4.9 (Fair Cut, [li2025simple, Theorem 1]).
@@
Consider the following flow problem: connect a source $s$ to all split nodes in $x_{a}\in X_{A}$ with capacity $\mu(x_{a})$ and connect all split nodes $x_{b}\in X_{B}$ to a target $t$ with capacity $\mu(x_{b})$. Using Theorem 4.9, we get a $2$-fair $(s,t)$-cut $(S,\bar{S})$ in $G^{\prime}$ corresponding to a $2$-approximate max flow $f$. We define $X_{Y}$ as follows: for every cut edge $(v,x_{e})\in E^{\prime}(S,\bar{S})$ we add $x_{e}$ to $X_{Y}$. This $X_{Y}$ separates $X_{A}$ from $X_{B}$. From the definition of fair-cuts and the fact that $\deg_{G^{\prime}}(x_{e})=2$, for every $x_{e}\in X_{Y}$, it follows that each $x_{e}\in X_{Y}$ receives (and also sends) at least $\operatorname{cap}(e)/2$ flow. Thus, by scaling $f$ by a factor of $2$ we get the desired outcome.
To see why $X_{Y}$ satisfies the constraints, decompose $f$ into paths. Since $(S,\bar{S})$ is a fair cut, each path crosses the cut exactly once (no flow can cross the cut backwards) and therefore can be assigned to a split vertex in $X_{Y}$. Thus, we can view $f$ as two routings: one from $s$ (i.e. $X_{A}$) to $X_{Y}$, and one from $X_{Y}$ to $t$ (i.e. $X_{B}$). Reversing the former flow gives the result. ∎
We define the second partition ($L$, $R$) by a computing the subset $X_{Y}$ of split nodes in $G^{\prime}[S^{\prime}]$, as obtained from applying Lemma 4.6 on $X_{F}$ and $X_{B}$ (see Figure 2). The sets $L$ and $R$ are then defined as follows: $R$ consists of all vertices in $S$ that can reach a boundary node $x_{b}\in X_{B}\setminus X_{Y}$ in the graph $G^{\prime}[S^{\prime}]\setminus X_{Y}$, while $L$ contains the remaining vertices. We adopt the notation of [racke2014computing] and call $L$ the core of $S$. Note that $X_{F}\setminus X_{Y}\subseteq L^{\prime}$, where $L^{\prime}\mathrel{\mathop{:}}=L\cup\{x_{e}\mid e=(u,v),u,v\in L\}$.
@@ 5 The Charging Scheme
Let $T$ be the tree corresponding to the hierarchical decomposition from Section 4. In this section we show that $T$ is indeed a tree cut-sparsifier with quality $O(\log^{3}n)$, proving Theorem 2.9. We define the capacity in $T$ of an edge connecting a cluster $S$ to its parent cluster $P(S)$, as $\operatorname{cap}_{G}(S,V\setminus S)$. In section 2 we mentioned that the method used to show that a tree is a cut-sparsifier is to prove that every multi-commodity flow problem that can be routed in the tree is respected by the graph. We begin by explaining this formally.
@@
If every multi-commodity flow problem (recall Definition 2.1) $R$ on $V$, that can be routed in $T$ with congestion $1$, can also be routed in $G$ with congestion $\alpha$, then $T$ is a tree flow-sparsifier with quality $\alpha$.
@@
Let $Q$ be a multi-commodity flow problem on $V$. We have to show that $\operatorname{cong}_{T}(Q)\leq\operatorname{cong}_{G}(Q)\leq\alpha\cdot\operatorname{cong}_{T}(Q)$. The first inequality is clear: We route the flow in $T$ naturally (a unit from $u$ to $v$ moves along the unique path connecting them). Therefore, there exists an edge $(S,P(S))$ of $T$ (of capacity $\operatorname{cap}_{G}(S,V\setminus S)$) such that $\operatorname{dem}_{Q}(S,V\setminus S)=\operatorname{cong}_{T}(Q)\cdot\operatorname{cap}_{G}(S,V\setminus S)$. Moreover, since $Q$ can be routed in $G$ with congestion $\operatorname{cong}_{G}(Q)$, it follows from Fact 2.7 that $\operatorname{dem}_{Q}(S,V\setminus S)\leq\operatorname{cong}_{G}(Q)\cdot\operatorname{cap}_{G}(S,V\setminus S)$. Therefore, we conclude that $\operatorname{cong}_{T}(Q)\leq\operatorname{cong}_{G}(Q)$.
For the other inequality, let $R$ denote the multi-commodity flow problem corresponding to $\frac{1}{\operatorname{cong}_{T}(Q)}\cdot Q$. This flow can be routed in $T$ with congestion $1$. Hence,
@@
The following is the corresponding claim for tree cut-sparsifiers. This will be our main tool to show that $T$ is a tree cut-sparsifier.
@@
If for every multi-commodity flow problem $R$ on $V$, that is $1$-respected by $T$, we have that $G$ $\frac{1}{\alpha}$-respects $R$ (recall Definition 2.6) then $T$ is a tree cut-sparsifier with quality $\alpha$.
@@
Fix a cut $(B,W)$. We have to show that $\operatorname{cap}_{G}(B,W)\leq\operatorname{mincut}_{T}(B,W)\leq\alpha\cdot\operatorname{cap}_{G}(B,W)$. The first inequality is clear given the definition of the capacities in $T$. Indeed, consider any cut $(B^{\prime},W^{\prime})$ in $T$, where $B\subseteq B^{\prime},W\subseteq W^{\prime}$. For each $b\in B,w\in W$, the path from the leaf $\{b\}$ to $\{w\}$ in $T$ must contain an edge $(S,P(S))$ that crosses the cut $(B^{\prime},W^{\prime})$ of $T$. Assume without loss of generality that $S\in B^{\prime}$ and $P(S)\in W^{\prime}$. Then, $\operatorname{cap}_{G}(b,w)$ will be counted in the capacity $\operatorname{cap}_{T}(S,P(S))=\operatorname{cap}_{G}(S,V\setminus S)$.
Next, we will show the other inequality. Consider the flow network $T^{\prime}=T\cup\{s,t\}$, where $s$ is connected to all nodes in $W$ and $t$ is connected to all nodes in $B$ with edges of infinite capacity. The maximum $(s,t)$-flow $f$ in this network has value $\operatorname{mincut}_{T}(B,W)$. Separate this flow into paths, and denote by $R$ the flow problem corresponding to the resulting flow demands (each demand has a distinct commodity, whose source is a node $b\in B$ and whose target is a node $w\in W$). Clearly, the congestion of routing $R$ in $T$ is $1$, which means (see Fact 2.7) that $R$ is $1$-respected by $T$. Therefore, we have $\frac{\operatorname{cap}_{G}(B,W)}{\operatorname{dem}_{R}(B,W)}\geq\frac{1}{\alpha}$. Note that $\operatorname{dem}_{R}(B,W)$, the total demand that crosses the cut, is exactly the value of the flow $|f|=\operatorname{mincut}_{T}(B,W)$. Therefore, we have $\operatorname{mincut}_{T}(B,W)\leq\alpha\cdot\operatorname{cap}_{G}(B,W)$, as we wanted. ∎
We will prefer to work with demand states instead of multi-commodity flow problems. To this end, we’ll use the following Corollary:
@@
Assume every valid demand state $P$ (recall Definition 3.14) on $V$, that is $1$-respected by $T$, is $\frac{1}{\alpha}$-respected by $G$. In other words, we have for any cut $(B,W)$,
$\frac{\operatorname{cap}_{G}(B,W)}{\operatorname{dem}_{P}(B,W)}\geq\frac{1}{\alpha}~.$
@@
Therefore, we have to show that every valid demand state that is respected by $T$ is $O(\log^{3}n)$-respected by $G$. To this end, we use $T$ as a blueprint that guides how demands are moved (in $G$) and charged to the capacity of cuts. Specifically, we update the demand state on $G$ iteratively as we go up along the tree structure, such that after processing a sub-cluster $S$, the demand state is no longer supported on any vertex of $S$, apart from it’s boundary split nodes. [^8] As will be explained soon, we work in the subdivision graph. As mentioned in Section 2, we will “pay” not for the congestion of routing the demands, but rather according to how much the graph respects the corresponding demand transitions.
The construction below follows [racke2014computing], with the required modifications for the case of cuts instead of flows. In contrast to [racke2014computing], our presentation makes use of demand states, which provides a convenient and precise framework for describing the dynamic movements of the demands as we go up along the decomposition tree and for analyzing how the graph respects them.
Fix a cut $(B\subseteq V,W=V\setminus B)$ in $G$, and fix a demand state $P$ on $V$, which is $1$-respected by $T$. Our goal is to show that $\operatorname{dem}_{P}(B,W)\leq O(\log^{3}n)\cdot\operatorname{cap}_{G}(B,W)$. We do this by charging the demand crossing the cut to the cut edges, such that the total charges dominate the demand, and each edge is charged $O(\log^{3}n)$.
@@
Let $\left(B^{\prime}=B\cup\{x_{e}\mid e\cap B\geq 1\},W^{\prime}=V^{\prime}\setminus B^{\prime}\right)$ be the corresponding cut in $G^{\prime}$ (see Figure 3). [^9] Note that we arbitrarily chose to include the split vertices of the edges crossing $(B,W)$ in $B^{\prime}$. Intuitively, we want to convert the demand state $P$ into a demand state $P^{\prime}$ on $G^{\prime}$, and then charge the edges crossing $(B^{\prime},W^{\prime})$ to cover the demand $\operatorname{dem}_{P^{\prime}}(B^{\prime},W^{\prime})$, which would be bounded from below by $\operatorname{dem}_{P}(B,W)$.
@@
Figure 3: Figure (a) - a cut $(B,W)$ in $G$. Figure (b) - the corresponding cut $(B^{\prime},W^{\prime})$ in $G^{\prime}$.
@@
[racke2014computing] essentially used Claim 5.1 to prove that $T$ is a tree flow-sparsifier, and to this end they showed how to route the corresponding valid demand state in $G$ with low congestion. Though they did not explicitly use the notation of demand states, their key method (rephrased in this notation) was akin to maintaining a collection of demand states, each supported on some sub-cluster $S^{\prime}$ of the subdivision graph, corresponding to a vertex of $T$. The original state $P$ was converted to a demand state $P_{\{v\}^{\prime}}$ supported on $\{v\}^{\prime}$ [^10] The notation $\{v\}^{\prime}$ refers to the cluster $S^{\prime}$ induced by the single vertex set $S=\{v\}$. Precisely, this includes the vertex $v$ as well as it’s adjacent split nodes., for each vertex $v\in V$ (these are the leaves in the decomposition tree $T$). They then showed how to merge the demand states according to the structure of the decomposition tree.

### v2

@@ 4.1 Partition of $G[S]$
      Assume $\tilde{A}\cup C$ induces a balanced clustering. We set $\tilde{C}\mathrel{\mathop{:}}=C$, $\tilde{F}=\tilde{A}\cup\tilde{C}$, and claim that we have Case (2) of the lemma. As mentioned above, the existence of the flow is guaranteed by Theorem 3.10, and each node $v\in A^{\prime}$ receives at most $\mu(v)$ units of flow. Since $\mu=\mu_{X_{F}}$, each node $x_{a}\in X_{\tilde{A}}=X_{F}\cap A^{\prime}$ receives at most one unit of flow, and the rest receive zero.
      The fact that $A^{\prime}$ $\Omega(\phi)$-expands with respect to $\mu_{X_{F}}$ in $G^{\prime}$ means that $G^{\prime}$ is an $\Omega(\phi)$ expander with respect to $X_{F}\cap A^{\prime}=X_{\tilde{A}}$, which by definition means that $X_{\tilde{A}}$ $\Omega(\phi)$-expands in $G^{\prime}$. As $\phi=\Theta(\frac{1}{\log n})$, this means that $X_{\tilde{A}}$ $\Omega(\frac{1}{\log n})$-expands in $G^{\prime}$. This shows we have Case (2) of the Lemma.
@@ 4.2 Finding a Bottleneck Cut
As mentioned in Section 2, we want the set $F\cup B$ of inter-cluster edges and boundary edges to expand in $G[S]$. However, the existence of bottleneck-cuts may make such a partition impossible (in the cut-sparsifier case, a bottleneck cut is a cut that certifies that the boundary edges are not expanding within the cluster). Similar to the technique [27] used in the flow-sparsifier case, the approach we use in this case is the following: As Claim 5.2 shows, in order to certify that a tree is a cut-sparsifier, we should check that all demand matrices which can be routed in the tree are respected by $G$. We therefore find a “problematic” cut $(L,R=S\setminus L)$ and add $L,R$ to the tree. This forces every demand matrix which is respected by $T$ to respect the cut $(L,V\setminus L)$. Specifically, we will add a cut which “blocks” the sending of flow from $X_{F}$ to $X_{B}$. As shown in Section 5, this will be enough.
Formally, the second partition is parameterized by $0<\uptau\leq 1$. The parameter $\uptau$ balances two opposing effects in our analysis: a larger $\uptau$ reduces the cost incurred when routing flow within a cluster, but simultaneously increases the $\ell_{1}$-blowup (that is, the increase in load) of the demand state vectors as we move from a cluster to its parent. Thus, $\uptau$ governs the trade-off between the internal routing cost and the load accumulation along the decomposition tree. See Theorem 5.5 for the exact details of this trade-off.
We shift our focus from the graph $G[S]^{\prime}$ to the graph $G^{\prime}[S^{\prime}]$, where $S^{\prime}\mathrel{\mathop{:}}=S\cup\{x_{e}\mid e\cap S\geq 1\}$. That is, in addition to the vertices of $G[S]^{\prime}$, we also include the nodes $x_{e}$ for every boundary edge $e\in B$ and connect $x_{e}$ to $v$ where $v\in e\cap S$.
Next, we define the weighting $\mu_{\uptau}$ on the vertices of $G^{\prime}[S^{\prime}]$: the boundary split nodes are assigned a weight of $\uptau$, and the rest of the vertices have weight 1.  [^9] For non split nodes, $v\in S$, $\mu_{\uptau}(v)$ does not play a role in the following discussion. As explained above, we are looking for a cut which blocks the sending of flow from $X_{F}$ to $X_{B}$. We therefore apply the following lemma on the sets $X_{F}$ and $X_{B}$. The set of edges $Y$ will be the cut edges that “block” the flow.
Lemma 4.5 (Improved version of [27, Lemma 3.9]).
@@
Consider the flow network where we add to $G^{\prime}$ a source $s$ connected to $X_{A}$ (with edges $(s,x_{a})$ of capacity $\mu(x_{a})$) and a sink $t$ connected to $X_{B}$ (with edges $(x_{b},t)$ of capacity $\mu(x_{b})$). It can be seen that an exact min–cut $Y$ that separates $s$ from $t$ in this network would satisfy the requirements of Lemma 4.5 with the constant $1$ instead of $2$. However, computing a min–cut is too expensive for our purposes. Instead, we use fair–cuts [21], which provide the same guarantees up to constant factors and can be computed in near-linear time.
Definition 4.7 (Fair Cut, [21]).
@@
Standard approximate max-flow algorithms have the unpleasant property that the flow provided can go backwards from $T$ to $S$ across the approximate min-cut $(S,T)$. Fair-cuts give a much stronger guarantee: a fair-cut requires that every edge that crosses the cut is roughly saturated. Thus, the flow can only cross the cut from $S$ to $T$ and the capacity of every cut edge is almost fully utilized. Intuitively, for our use cases, fair-cuts can be thought of as analogues to minimum cuts which can be computed in near linear time (they give the same guarantees as minimum cuts, up to constant factors).
The following theorem, due to [20], is the state of the art result for the computation of fair-cuts.
Theorem 4.8 (Fair Cut, [20, Theorem 1]).
@@
Consider the following flow problem: start with $G^{\prime}$ and set the edge capacities according to $\mu$. That is, set $\operatorname{cap}(u,x_{e})=\operatorname{cap}(x_{e},v)=\mu(x_{e})$, for every $e=(u,v)\in E$. Next, connect a source $s$ to all split nodes in $x_{a}\in X_{A}$ with capacity $\mu(x_{a})$ and connect all split nodes $x_{b}\in X_{B}$ to a target $t$ with capacity $\mu(x_{b})$. Using Theorem 4.8, we get a $\frac{3}{2}$-fair $(s,t)$-cut $(S,\bar{S})$ in $G^{\prime}$ corresponding to a $\frac{3}{2}$-approximate max flow $f$. We define $X_{Y}$ as follows: for every cut edge $(v,x_{e})\in E^{\prime}(S,\bar{S})$ we add $x_{e}$ to $X_{Y}$. This $X_{Y}$ separates $X_{A}$ from $X_{B}$. From the definition of a $\frac{3}{2}$-fair cut and since $X_{Y}$ consists of split nodes, it follows that each $x_{e}\in X_{Y}$ receives (and also sends) at least $\frac{2}{3}\mu(x_{e})$ flow.
The desired flows are obtained by the following modification of $f$. Consider the scaled flow $f^{\prime}:=\frac{3}{2}f$. For each $x_{e}\in X_{Y}$, the total flow (in $f^{\prime}$) $x_{e}$ sends (or receives, depends on its side in the cut) through the cut $(S,\bar{S})$ is at least $\mu(x_{e})$. We decompose $f^{\prime}$ into flow paths and scale them down such that, for $x_{e}\in X_{Y}$, the total flow $x_{e}\in X_{Y}$ sends (or receives) through $(S,\bar{S})$ is exactly $\mu(x_{e})$. This scaling argument is possible since $(S,\bar{S})$ is a fair-cut and therefore each flow path of $f$ crosses the cut $(S,\bar{S})$ exactly once. Let $f^{\prime\prime}$ be the resulting flow.
To see why $X_{Y}$ satisfies the constraints, decompose $f^{\prime\prime}$ into flow paths. Thus, we can view $f^{\prime\prime}$ as two routings: one from $s$ (i.e., $X_{A}$) to $X_{Y}$, and one from $X_{Y}$ to $t$ (i.e., $X_{B}$). Reversing the former flow gives the result. ∎
We define the second partition ($L$, $R$) by computing the subset $X_{Y}$ of split nodes in $G^{\prime}[S^{\prime}]$, as obtained from applying Lemma 4.5 on $X_{F}$ and $X_{B}$ (see Figure 2). The sets $L$ and $R$ are then defined as follows: $R$ consists of all vertices in $S$ that can reach a boundary node $x_{b}\in X_{B}\setminus X_{Y}$ in the graph $G^{\prime}[S^{\prime}]\setminus X_{Y}$, while $L$ contains the remaining vertices. We adopt the notation of [27] and call $L$ the core of $S$. Note that $X_{F}\setminus X_{Y}\subseteq L^{\prime}$, where $L^{\prime}\mathrel{\mathop{:}}=L\cup\{x_{e}\mid e=(u,v),u,v\in L\}$.
@@
Remark 4.10.
In degenerate cases, it may occur that $F=\emptyset$ (e.g., when $G[S]$ has no edges) or $B=\emptyset$ (when $S=V$ or if $G$ is disconnected). In either case, Theorem 4.9 returns $Y=\emptyset$.
At this point, we can give some informal intuition for the role of the parameter $\uptau$ (used in Theorem 4.9) and the trade-off it induces. As we will see in the next section, to prove that our hierarchical decomposition $T$ is a tree cut-sparsifier, it suffices to show that at each cluster $S$ we can perform the following local operation: given demands supported on the boundaries of the sub-clusters of $S$ (that is, on $X_{F}\cup X_{Y}\cup X_{B}$), we are able to (i) sufficiently mix the demands inside $S$, so that some positive and negative demands on $X_{F}\cup X_{Y}$ cancel out (see Lemma 5.9 and Corollary 5.11), and (ii) route the remaining demands to the boundary split nodes $X_{B}$ of $S$, while ensuring that the load on each boundary split node does not increase too much, so that these demands can be handled at higher levels of the tree. Theorem 4.9 provides the mechanism need to carry these operations.
For operation (i), in order to mix the demands on $X_{F}\cup X_{Y}$, we first move the demands residing on $X_{Y}$ to $X_{F}$, and then exploit the fact that $F$ $\Omega(\frac{1}{\log n})$-expands in $S$ in order to mix these demands inside $X_{F}$. We therefore apply Theorem 4.9 and move the demands from $X_{Y}$ to $X_{F}$. Since $X_{Y}$ may include boundary split nodes $x_{y}\in X_{Y}\cap X_{B}$ (which have constant load on them but a capacity of $\mu_{\uptau}(x_{y})=\uptau$), the congestion incurred by routing the demands from $X_{Y}$ to $X_{F}$ is inversely proportional to $\uptau$.
For operation (ii), we use the fact that the cut $(L,R)$, constructed using the separator $Y$, was added to the hierarchical decomposition. This allows us to argue that after the mixing performed in operation (i), the total demand remaining on $X_{F}\cup X_{Y}$ is bounded by the capacity of YY [^10] Actually, the demand is bounded by the capacity of $\tilde{Y}\subseteq Y$ which is defined in the proof. (see Corollary 5.11). Consequently, by again invoking the routing guarantees of Theorem 4.9, we can route the remaining demands to $X_{Y}$ such that the total demand on each vertex is constant. Finally, by applying Theorem 4.9 once more, we move the demands from $X_{Y}$ to the boundary $X_{B}$, while increasing the load on the boundary split nodes by at most a factor of $1+\uptau$.
@@ 5 The Charging Scheme
Let $T$ be the tree corresponding to the hierarchical decomposition from Section 4. In this section we show that $T$ is indeed a tree cut-sparsifier with quality $O(\log^{3}n)$, proving Theorem 2.9. We define the capacity in $T$ of an edge connecting a cluster $S$ to its parent cluster $P_{T}(S)$, as $\operatorname{cap}_{G}(S,V\setminus S)$. In section 2 we mentioned that the method used to show that a tree is a cut-sparsifier is to prove that every demand matrix that can be routed in the tree is respected by the graph. We begin by explaining this formally.
@@
The following notions are equivalent for a (hierarchical) tree $T$ and a parameter $\alpha$:
  • $T$ is a tree flow-sparsifier of $G$ with quality $\alpha$.
  • For every demand matrix (recall Definition 2.1) $Q$ on $V$, if $Q$ can be routed in $T$ with congestion $1$, then it can be routed in $G$ with congestion $\alpha$.
@@
The implication from the first item to the second is immediate from the definition of a flow-sparsifier. We therefore prove the converse.
Let $Q$ be a demand matrix on $V$. We have to show that $\operatorname{cong}_{T}(Q)\leq\operatorname{cong}_{G}(Q)\leq\alpha\cdot\operatorname{cong}_{T}(Q)$. The first inequality is clear: We route the flow in $T$ naturally (a unit from $u$ to $v$ moves along the unique path connecting them). Therefore, there exists an edge $(S,P_{T}(S))$ of $T$ (of capacity $\operatorname{cap}_{G}(S,V\setminus S)$) such that $\operatorname{dem}_{Q}(S,V\setminus S)=\operatorname{cong}_{T}(Q)\cdot\operatorname{cap}_{G}(S,V\setminus S)$. Moreover, since $Q$ can be routed in $G$ with congestion $\operatorname{cong}_{G}(Q)$, it follows from Fact 2.7 that $\operatorname{dem}_{Q}(S,V\setminus S)\leq\operatorname{cong}_{G}(Q)\cdot\operatorname{cap}_{G}(S,V\setminus S)$. Therefore, we conclude that $\operatorname{cong}_{T}(Q)\leq\operatorname{cong}_{G}(Q)$.
For the other inequality, let $R$ denote the flow problem corresponding to $\frac{1}{\operatorname{cong}_{T}(Q)}\cdot Q$. This flow can be routed in $T$ with congestion $1$. Hence,
@@
The following is the corresponding claim for tree cut-sparsifiers. It is our main tool to show that $T$ is a tree cut-sparsifier. Recall the definition of “respecting a cut” (Definition 2.6).
@@
The following notions are equivalent for a (hierarchical) tree $T$ and a parameter $\alpha$:
  • $T$ is a tree cut-sparsifier of $G$ with quality $\alpha$.
  • For every demand matrix $Q$ on $V$, if $T$ 11-respects $Q$, then $G$ $\frac{1}{\alpha}$-respects $Q$.
@@
The implication from the first item to the second is immediate from the definition of a cut-sparsifier. We therefore prove the converse.
Fix a cut $(U,W)$ of $V$. We have to show that $\operatorname{cap}_{G}(U,W)\leq\operatorname{mincut}_{T}(U,W)\leq\alpha\cdot\operatorname{cap}_{G}(U,W)$. The first inequality is clear given the definition of the capacities in $T$. Indeed, consider any cut $(U^{\prime},W^{\prime})$ in $T$, where $U\subseteq U^{\prime},W\subseteq W^{\prime}$. For each $u\in U,w\in W$, the path from the leaf $\{u\}$ to $\{w\}$ in $T$ must contain an edge $(S,P_{T}(S))$ that crosses the cut $(U^{\prime},W^{\prime})$ of $T$. Assume without loss of generality that $S\in U^{\prime}$ and $P_{T}(S)\in W^{\prime}$. In particular, $u\in S$. Then, $\operatorname{cap}_{G}(u,w)$ will be counted in the capacity $\operatorname{cap}_{T}(S,P_{T}(S))=\operatorname{cap}_{G}(S,V\setminus S)$.
Next, we will show the other inequality. Consider the flow network $T^{\prime}=T\cup\{s,t\}$, where $s$ is connected to all nodes in $W$ and $t$ is connected to all nodes in $U$ with edges of infinite capacity. Due to the max-flow min-cut theorem, the maximum $(s,t)$-flow $f$ in this network has value $\operatorname{mincut}_{T}(U,W)$. Separate this flow into paths. Each path begins in a vertex of $U$ and ends in a vertex of $W$. Denote by $R$ the demand matrix corresponding to the flow paths (each demand has a distinct commodity, whose source is a node $u\in U$ and whose target is a node $w\in W$). Clearly, the congestion of routing $R$ in $T$ is $1$, which means (see Fact 2.7) that $R$ is $1$-respected by $T$. Therefore, we have $\frac{\operatorname{cap}_{G}(U,W)}{\operatorname{dem}_{R}(U,W)}\geq\frac{1}{\alpha}$. Note that $\operatorname{dem}_{R}(U,W)$, the total demand that crosses the cut, is exactly the value of the flow $|f|=\operatorname{mincut}_{T}(U,W)$. Therefore, we have $\operatorname{mincut}_{T}(U,W)\leq\alpha\cdot\operatorname{cap}_{G}(U,W)$, as we wanted. ∎
We prefer to work with demand states instead of demand matrices. To this end, we use the following Corollary:
@@
Assume every valid demand state $P$ (recall Definition 3.13) on $V$ (with arbitrary commodities $\mathcal{K}$), that is $1$-respected by $T$, is $\frac{1}{\alpha}$-respected by $G$. In other words, we have for any cut $(U,W)$,
$\frac{\operatorname{cap}_{G}(U,W)}{\operatorname{dem}_{P}(U,W)}\geq\frac{1}{\alpha}~.$
@@
Therefore, we have to show that every valid demand state that is respected by $T$ is $\Omega\!\left(\frac{1}{\log^{3}n}\right)$-respected by $G$. To this end, we use $T$ as a blueprint that guides how demands are moved (in $G$) and charged to the capacity of cuts. Specifically, we update the demand state on $G$ iteratively as we go up along the tree structure, such that after processing a sub-cluster $S$, the demand state is no longer supported on any vertex of $S$, apart from it’s boundary split nodes. [^11] As will be explained soon, we work in the subdivision graph. As mentioned in Section 2, we will “pay” not for the congestion of routing the demands, but rather according to how much the graph respects the corresponding demand transitions.
The construction below follows [27], with the required modifications for the case of cuts instead of flows. In contrast to [27], our presentation makes use of demand states, which provides a convenient and precise framework for describing the dynamic movements of the demands as we go up along the decomposition tree and for analyzing how the graph respects them.
Fix a cut $(U\subseteq V,W=V\setminus U)$ in $G$, and fix a valid demand state $P$ on the vertex set $V$ (with arbitrary commodity set $\mathcal{K}$), which is $1$-respected by $T$. Our goal is to show that $\operatorname{dem}_{P}(U,W)\leq O(\log^{3}n)\cdot\operatorname{cap}_{G}(U,W)$. We do this by charging the demand crossing the cut to the cut edges, such that the total charges dominate the demand, and each edge is charged $O(\log^{3}n)$.
@@
Let $\left(U^{\prime}=U\cup\{x_{e}\mid e\cap U\geq 1\},W^{\prime}=V^{\prime}\setminus U^{\prime}\right)$ be the corresponding cut in $G^{\prime}$ (see Figure 3). [^12] Note that we arbitrarily chose to include the split vertices of the edges crossing $(U,W)$ in $U^{\prime}$. Intuitively, we want to convert the demand state $P$ into a demand state $P^{\prime}$ on $G^{\prime}$, and then charge the edges crossing $(U^{\prime},W^{\prime})$ to cover the demand $\operatorname{dem}_{P^{\prime}}(U^{\prime},W^{\prime})$, which would be bounded from below by $\operatorname{dem}_{P}(U,W)$.
@@
Figure 3: Figure (a) - a cut $(U,W)$ in $G$. Figure (b) - the corresponding cut $(U^{\prime},W^{\prime})$ in $G^{\prime}$.
@@
[27] essentially used Claim 5.1 to prove that $T$ is a tree flow-sparsifier, and to this end they showed how to route the corresponding valid demand state in $G$ with low congestion. Though they did not explicitly use the notation of demand states, their key method (rephrased in this notation) was akin to maintaining a collection of demand states, each supported on some sub-cluster $S^{\prime}$ of the subdivision graph, corresponding to a vertex of $T$. The original state $P$ is converted to a set of demand states $\{P_{\{v\}^{\prime}}\}_{v\in V}$, where for each vertex $v\in V$, $P_{\{v\}^{\prime}}$ is supported on $\{v\}^{\prime}$ [^13] The notation $\{v\}^{\prime}$ refers to the cluster $S^{\prime}$ induced by the single vertex set $S=\{v\}$. That is, this includes the vertex $v$ as well as it’s adjacent split nodes. (the vertices $V$ are the leaves in the decomposition tree $T$). Then they show how to merge the demand states according to the structure of the decomposition tree.